# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 3 Estimates for Discrete Green’s function
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}N\leq\frac{C}{N}.$ (3.36)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{7/2}}{N^{3/2}}.$
@@
$\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Lemma 4.4 When $\alpha>2$, $\beta>2/3$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}$.
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})$ $=\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{c}}$ (3.37)
Now, consider the first term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\quad\frac{\varepsilon^{1/2}}{\varepsilon^{2}}\sqrt{|\Omega_{f,xy}|}\;N\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon^{3/2})^{2}}{\varepsilon N}\varepsilon^{3/4}\log(1/\varepsilon^{3/2})$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N}.$ (3.38)
Now, consider the first term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,xy}}$
@@
$\leq C\varepsilon^{1/2}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\frac{\varepsilon^{3/4}}{\varepsilon^{2}}\log(1/\varepsilon^{3/2})N$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.39)
Now, consider second term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\sqrt{|\Omega_{f,x}|}\;N\leq C\frac{(\varepsilon\log(1/\varepsilon))^{1/2}}{\varepsilon^{3/2}N}$
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}.$ (3.40)
$\text{Note that}\quad\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}=1.$ (3.41)
Consider $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}.$
Note that when $(x,y)\in\Omega_{f,x}$, then $x\in[0,\varepsilon\log(1/\varepsilon)]$ and $y\in[-1+\lambda_{y},1-\lambda_{y}]$.
If $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}=\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$.
Let $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}=e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y_{0})}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y_{0})}$ where $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$.
Let $y_{0}=-1+\lambda_{y}$ then $1+y_{0}=\lambda_{y}=\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)$
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y_{0})}$ $=e^{-\frac{\beta}{\sqrt{\varepsilon}}\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{-\beta\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\frac{3\beta}{2}}}\right)}=\varepsilon^{\frac{3\beta}{2}}.$ (3.42)
and $1-y_{0}=1-(-1+\lambda_{y})=2-\lambda_{y}=2-\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)\geq 1$ for sufficiently small $\varepsilon>0$.
$e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y_{0})}$ $=\frac{1}{e^{\frac{\beta}{\sqrt{\varepsilon}}(1-y_{0})}}=\frac{1}{1+\frac{\beta}{1!\sqrt{\varepsilon}}+\frac{\beta^{2}}{2!\sqrt{\varepsilon}^{2}}+\frac{\beta^{3}}{3!\sqrt{\varepsilon}^{3}}+\dots}\leq C\varepsilon^{\frac{3}{2}\beta}.$ (3.43)
Combining (3.40), (3.41), (3.42) and (3.43), we obtain
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq C\frac{(\log(1/\varepsilon))^{1/2}}{\varepsilon N}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}$
$\leq C\frac{(\log(1/\varepsilon))^{1/2}}{\varepsilon N}\varepsilon^{\frac{3}{2}\beta}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.\quad\text{when}\;\beta>2/3.$ (3.44)
Now, consider second term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,x}}$
@@
$\quad\sqrt{|\Omega_{f,x}|}\;N\leq C\frac{1}{N}\varepsilon^{1/2}(\log(1/\varepsilon))^{5/2}\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N}.$ (3.45)
Now, consider the third term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\quad\frac{\varepsilon^{1/2}}{\varepsilon^{2}}\sqrt{|\Omega_{f,y}|}\;N\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{\varepsilon^{1/4}N}e^{\alpha}$
$\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\quad\text{when}\;\alpha>1/4.$ (3.46)
Now, consider the third term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,y}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{5/2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;N$
$\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{\varepsilon^{7/4}N}e^{\alpha}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{N}\quad\text{when}\;\alpha>7/4.$ (3.47)
Now, consider the fourth term of the R.H.S. of in the $x$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;N$
$\leq C\frac{1}{\varepsilon^{3/2}N}e^{\alpha}\leq\frac{C}{N}\quad\text{when}\;\alpha>3/2.$ (3.48)
Now, Consider the fourth term of the R.H.S. of in the $y$-direction (3.37) $\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{c}}$. We obtain,
$\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{c}}$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{c}}$
@@
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{5/2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;N$
$\leq C\frac{1}{\varepsilon^{2}N}e^{\alpha}\leq\frac{C}{N}\quad\text{when}\;\alpha>2.$ (3.49)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{5/2}}{N^{3/2}}.$
@@
$\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
4 Estimates for convection and reaction terms involving discrete Green’s function
In this section, we aim to bound the convection and reaction portion of the bilinear form B(.,.) with its arguments parts of interpolation error uIuu-Iu and the discrete Green’s function $G_{mn}.$ Finally, based on the estimates for the diffusion, convection, and reaction terms, we derive the main theorem of this article.
Theorem 3

### v2

@@ 4 Estimates related to diffusion terms
$\leq C\varepsilon^{1/2}\frac{1}{N^{2}}\frac{1}{\varepsilon^{1/2}}\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}.$ (4.26)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
@@
$|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Lemma 4.5
If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then for $\alpha\geq 2,\;\beta\geq 2/3$ following estimate hold $|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{f,x}$ and decompose the domain in four parts, we have
$|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})|$ $\leq|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,xy}}|+|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,x}}|$
$+|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{f,y}}|+|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})_{\Omega_{c}}|$ (4.27)
Now, estimate first term of the R.H.S. of (4.27) in the $x$-direction is $|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,xy}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\frac{\varepsilon^{1/2}}{\varepsilon^{2}}\sqrt{|\Omega_{f,xy}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{5/4}\frac{(\log\left(1/\varepsilon^{3/2}\right))^{4}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{4}}{N^{3/2}}.$ (4.28)
Now, estimate first term of the R.H.S. of (4.27) in the $y$-direction is $|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,xy}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,xy}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,xy}}$
@@
$\leq C\varepsilon^{1/2}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\frac{\varepsilon^{3/4}}{\varepsilon^{2}}\log(1/\varepsilon^{3/2})\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}.$ (4.29)
Now, estimate second term of the R.H.S. of (4.27) in the $x$-direction is $|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Lemma 3.4 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,x}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq\frac{C\varepsilon^{1/2}}{\varepsilon^{2}N^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}\;\|G_{mn}\|_{{{}_{1,\varepsilon}}}$
$\leq\frac{C(\varepsilon\log(\frac{1}{\varepsilon}))^{1/2}}{\varepsilon^{3/2}N}\varepsilon^{\frac{3}{2}\beta}\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\log(1/\varepsilon)}{N}\varepsilon^{\frac{3}{2}\beta-1}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}\quad\text{when}\;\beta\geq 2/3.$ (4.30)
Now, estimate second term of the R.H.S. of (4.27) in the $y$-direction is $|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Lemma 3.4 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,x}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,x}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,x}}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\sqrt{|\Omega_{f,x}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{1}{N^{3/2}}\varepsilon^{1/2}(\log(1/\varepsilon))^{3}\varepsilon^{\frac{3\beta}{2}}\leq C\frac{(\log(1/\varepsilon))^{3}}{N^{3/2}}.$ (4.31)
Now, estimate third term of the R.H.S. of (4.27) in the $x$-direction is $|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{f,y}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\frac{\varepsilon^{1/2}}{\varepsilon^{2}}\sqrt{|\Omega_{f,y}|}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{\varepsilon^{1/4}N^{3/2}}\varepsilon^{\alpha}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}\quad\text{when}\;\alpha\geq 1/4.$ (4.32)
Now, estimate third term of the R.H.S. of (4.27) in the $y$-direction is $|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3, Corollary 3.4.1 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{f,y}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{f,y}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{f,y}}$
@@
$\leq\frac{C\sqrt{\varepsilon}}{N^{2}\varepsilon^{\frac{5}{2}}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\sqrt{N\log\frac{1}{\varepsilon}}$
$\leq C\frac{\log(1/\varepsilon^{3/2})}{\varepsilon^{7/4}N^{3/2}}\varepsilon^{\alpha}\leq C\frac{\log(1/\varepsilon^{3/2})}{N^{3/2}}\quad\text{when}\;\alpha\geq 7/4.$ (4.33)
Now, estimate fourth term of the R.H.S. of (4.27) in the $x$-direction is $|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3, Lemma 3.4 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{x},G_{mn,x})_{\Omega_{c}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{x}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,x}\|_{\Omega_{c}}$
@@
$\frac{\leq C\varepsilon^{1/2}}{N^{2}}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\sqrt{N\log\frac{1}{\varepsilon}}$
$\leq\frac{C}{\varepsilon^{3/2}N^{3/2}}\sqrt{\log\frac{1}{\varepsilon}}\varepsilon^{\alpha+\frac{3\beta}{2}}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}\quad\text{when}\;\alpha+\frac{3\beta}{2}\geq 3/2.$ (4.34)
Now, estimate fourth term of the R.H.S. of (4.27) in the $y$-direction is $|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3, Lemma 3.4 and furthermore, we obtain
$|\varepsilon((F_{12}-IF_{12})_{y},G_{mn,y})_{\Omega_{c}}|$ $\leq\sqrt{\varepsilon}\|(F_{12}-IF_{12})_{y}\|_{\Omega_{c}}\sqrt{\varepsilon}\|G_{mn,y}\|_{\Omega_{c}}$
@@
$\leq C\frac{\varepsilon^{1/2}}{\varepsilon^{\frac{5}{2}}N^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\sqrt{N\log\frac{1}{\varepsilon}}$
$\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N^{3/2}}\varepsilon^{\alpha+\frac{3\beta}{2}-2}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}\quad\text{when}\;\alpha+\frac{3\beta}{2}\geq 2.$ (4.35)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{7/2}}{N}.$
@@
$|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{7/2}}{N}.$
5 Estimates related to convection and reaction terms
In this section, we aim to bound the convection and reaction portion of the bilinear form B(.,.) with its arguments parts of interpolation error uIuu-Iu and the discrete Green’s function $G_{mn}.$ For the convection term, while we get $\varepsilon$-independent estimate for $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, the convergence behavior differs. For $(x_{m},y_{n})\in\Omega_{c}$, we could not determine a specific order of convergence. For $(x_{m},y_{n})\in\Omega_{f,x}$, we obtained a convergence order of $N^{-1/2}$ up to a logarithmic factor. And for the reaction term, we got a convergence order of $N^{-1}$ up to a logarithmic factor when $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$. We begin our section with the following auxiliary results.
Theorem 5.1